Two cities are 315 miles apart. a car leaves one of the cities traveling toward the second city at 50 miles per hour. At the same time a bus leaves the second city at 55 miles per hour. How long will it take for them to meet?
Answer by edjones(8007)

Since they are 315 mi apart and the are traveling toward each other their commbined distance=315 mi.
distance=speed*time
d=s*t
for the car d=50t
for the bus d=55t
50t+55t=315
105t=315
divide 105 into both sides. t=3 hrs
Check
50*3+55*3=150+165=315
Ed